# Safe and Optimal N-Spacecraft Swarm Reconfiguration in Non-Keplerian Cislunar Orbits

### v1

@@ 3.2 Local Toroidal Coordinates
One can also define LTC by expressing the eigenspace component in polar coordinates (cf. Eq. (11)) as $\boldsymbol{\zeta}_{\text{geo}}:=[\varepsilon,\theta,h,\dot{\varepsilon},\dot{\theta},\dot{h}]^{\top}$, which is referred to as geometric LTC. This approximation holds for the relative motion around any periodic orbit thanks to the linearization scheme adopted in Ref. 16, and the validity of its accuracy is directly affected by the size and direction (eigenspace) of the invariant circle at each phase of the reference orbit. A similar mapping from the geometric LTC to Cartesian coordinates is available [22], although the transformation is no longer a function of only the states on a reference orbit but also a time derivative of nonsingular LTC. Nonetheless, the geometric LTC provides a convenient characterization of an invariant circle’s size, thereby facilitating the specification of boundary conditions and waypoints for trajectory design.
The first-order condition to be on a QPRIT (QPRO) is to stay on the eigenspace spanned by $\boldsymbol{w}_{r}(t)$ and $\boldsymbol{w}_{i}(t)$. This leads to the condition $h=\dot{\alpha}=\dot{\beta}=\dot{h}=0$, or using geometric LTC, $h=\dot{\varepsilon}=\dot{\theta}=\dot{h}=0$ [26]. Any other coordinates still define the deputy’s relative motion, although the (first-order) quasi-periodicity is no longer attainable. The quasi-periodic condition allows for arbitrary choices of the variables $(\alpha,\beta)$ or $(\varepsilon,\theta)$, defining both the size of the invariant circle and its phase. However, nonzero values in $(\dot{\alpha},\dot{\beta})$ or $(\dot{\varepsilon},\dot{\theta})$ result in a drift in the $h$-component, breaking the quasi-periodicity.
@@ 3.3 Safety evaluation of QROs
This subsection demonstrates the connection between the ROE and LTC based on the geometric configuration of the osculating relative orbit and the minimum separation. Although the approach is general, the application case of this section is on the Earth-Moon $L_{2}$ 9:2 South NRHO, which is an orbit of a Lunar Gateway [2, 41] and has an orbital period of 6.56 days.
3.3.1 Comparison between the LVLH and TNW frame
The first-order approximation of a QPRIT resolved in the LVLH and TNW frames is illustrated in Figs. 5(a) and 5(b), respectively. Both figures depict the same QPRIT, where the fixed point is anchored at apolune with a scaling factor $\varepsilon=200$ m. The surfaces of the QPRIT are generated by the continuation of invariant circles (cf. Eq. (11)) along the reference orbit, shown as a gray surface. Additionally, fifty osculating invariant circles, sampled uniformly over one orbital period, are shown as blue ellipses. In the TNW frame, the semimajor axes of the invariant circles align with the chief’s velocity direction, producing a more compact structure. In contrast, for highly eccentric orbits like NRHOs, the angular offset of the QPRIT in the LVLH frame is evident. Previous research [26, 22, 42] reports a QPRIT structure of 9:2 NRHO characterized by two disk-like surfaces, a pattern arising from positional displacement relative to the spacecraft velocity in the synodic frame. The reduced density of invariant circles within these disks reflects higher relative velocities and greater radial uncertainty.
@@ 3.3.2 Geometric Interpretation of Osculating Relative Invariant Circle
Two-dimensional projections of an osculating relative orbit in the PR2BP and an osculating invariant circle in the RMBPs are shown in Figs. 6(a) and 6(b), respectively. For a relative motion around a Keplerian orbit, a six-dimensional variable $\boldsymbol{\kappa}=[\kappa_{1},\kappa_{2},\kappa_{3},\kappa_{4},\kappa_{5},\kappa_{6}]\in\mathbb{R}^{6}$ parameterizes a relative orbit of the deputy with respect to the chief (cf. Eq. (9)), with additional independent variable $\rho(t)=1+e\cos\nu(t)$, where $e$ and $\nu(t)$ denote the eccentricity and true anomaly of the chief, respectively. In this paper, quasi-nonsingular ROE in near-circular orbits [5] and its extension to elliptic orbits [40] are chosen for $\boldsymbol{\kappa}$, which correspond to integration constants of the linearized equations of motion [3, 4]. In Fig. 6(a), $\kappa_{1}=0$ is enforced to realize a non-drifting bounded relative motion.
An osculating invariant circle in Eq. (11) can be parameterized as
$\boldsymbol{\psi}(t)=\varepsilon\begin{bmatrix}\sqrt{c_{1}^{2}(t)+c_{2}^{2}(t)}\sin(\theta+\phi_{x}(t))\\ \sqrt{c_{3}^{2}(t)+c_{4}^{2}(t)}\sin(\theta+\phi_{y}(t))\\ \sqrt{c_{5}^{2}(t)+c_{6}^{2}(t)}\sin(\theta+\phi_{z}(t))\end{bmatrix},\quad\boldsymbol{r}_{r}(t)=\begin{bmatrix}c_{1}(t)\\ c_{3}(t)\\ c_{5}(t)\end{bmatrix},\quad\boldsymbol{r}_{i}(t)=\begin{bmatrix}c_{2}(t)\\ c_{4}(t)\\ c_{6}(t)\end{bmatrix},\quad\begin{cases}\tan\phi_{x}(t)=c_{1}(t)/c_{2}(t)\\ \tan\phi_{y}(t)=c_{3}(t)/c_{4}(t)\\ \tan\phi_{z}(t)=c_{5}(t)/c_{6}(t)\end{cases}.$ (16)
@@
(b) Osculating invariant circle of a QPRO (RMBP) in the TNW frame based on the parametrization in Eq. (16).
@@
Figure 6: Comparison of the geometric parametrization of osculating relative orbit in the PR2BP and the invariant circle in the CR3BP.
@@
Vectors $[c_{1}(t),c_{2}(t)]$, $[c_{3}(t),c_{4}(t)]$, and $[c_{5}(t),c_{6}(t)]$ provide the amplitudes of the projected ellipse in each plane as well as the oscillation phase. As stated in Sec. 3.1.1, $\boldsymbol{\kappa}$ is time-invariant [^5] When perturbations (e.g., spherical harmonics, drag, etc.) are added to the Keplerian dynamics, the osculating ROE also experiences a drift, although it does not induce any quasi-periodic motion. Nonetheless, the magnitude of the perturbation on the eigensystem is negligible (i.e., the eigenvectors of the monodromy matrix are almost 1) and can be treated as a secular drift in an averaged motion over multiple revolutions in a phase space. in the R2BP, allowing the explicit expression of the minimum planar separation in each projected plane [5]. In contrast, vector $\boldsymbol{c}(t)=[c_{1}(t),c_{2}(t),c_{3}(t),c_{4}(t),c_{5}(t),c_{6}(t)]\in\mathbb{R}^{6}$ is time-varying in RMBPs, where a combined analysis of the magnitudes and phases of this vector is necessary to assess the passive safety of a QPRO.
@@
The history of $\boldsymbol{c}(t)$ (i.e., eigenvector $\boldsymbol{w}(t)$) over eight revolutions is shown in Fig. 7, resolved in the LVLH frame and the TNW frame, respectively. Note that this fixes the size of the osculating invariant circle $\varepsilon=1$, while the geometric structure is preserved. Fig. 7 is analogous to the plots of the vector fields of relative eccentricity $[\kappa_{3},\kappa_{4}]$ and inclination $[\kappa_{5},\kappa_{6}]$ vector in the quasi-nonsingular ROE [5, 40] because the magnitude of each vector represents the amplitude of osculating invariant circle in each direction of the axis ($c_{12},c_{34}$, and $c_{56}$), and the magnitudes of the two vectors and the phase difference between them determine the minimum separation in the corresponding two-dimensional projected plane. However, note again that the vectors $[c_{1}(t),c_{2}(t)]$, $[c_{3}(t),c_{4}(t)]$, and $[c_{5}(t),c_{6}(t)]$ for a unit invariant circle are determined by the local eigensystem of the chief’s absolute position and velocity, whereas the relative eccentricity and inclination vectors are free design variables of the deputy.
@@
From each vector field illustrated in Fig. 7, a quasi-periodic behavior is confirmed with a drastic change in its amplitude. Note that the equivalent vectors in relative motion for the PR2BP (i.e., ROE) would be stationary (for a near-circular orbit) or oscillatory about a line segment (for an elliptic orbit), with possible slow secular drifts due to perturbations. The evolution the eigenvector can be characterized by two angles $\boldsymbol{\theta}=[\theta_{0},\theta_{1}]\in\mathbb{T}^{2}$ by separating the magnitude and the phase as $\boldsymbol{w}(t)=w_{r}(\theta_{0}(t))\hat{\boldsymbol{w}}_{r}(\theta_{1}(t))+w_{i}(\theta_{0}(t))\hat{\boldsymbol{w}}_{i}(\theta_{1}(t))i$. The first angle, $\theta_{0}$, determines the magnitude of the vector corresponding to the fixed point on the reference orbit. In contrast, the second angle, $\theta_{1}$, captures the secular drift of the vector over multiple orbital revolutions, reflecting the geometric structure of the two-dimensional QPIT and QPRIT (cf. Fig. 3(b)). The angular rate of this secular rotation $\dot{\theta}_{1}$ remains constant over the three vector fields and corresponds to the angle of the complex eigenvalue of the oscillatory mode, known as rotation number [33, 34]. For example, an eigenvalue associated with the oscillatory mode of the monodromy matrix of the 9:2 South NRHO is $\lambda\simeq 0.6845-0.7290i$. This leads to the rotation number of $\varrho\simeq-46.80^{\circ}$, which coincides with the angular displacement between the initial and terminal points of the orange segment in Fig. 7. For both Figs. 7(a) and 7(b), the minimum amplitude in the N-direction collapses to nearly zero, as shown in the history of $[c_{3}(t),c_{4}(t)]$. A significant difference between Figs. 7(a) and 7(b) lies in the vector $[c_{5}(t),c_{6}(t)]$, which represents the amplitude of the invariant circle in the R(W)-direction. This is illustrated by the contraction of the amplitude in the W-direction in the TNW frame compared to the R-direction in the LVLH frame.
@@
(b) TNW frame
Figure 7: History of $\boldsymbol{r}_{r}(t)=[c_{1}(t),c_{3}(t),c_{5}(t)]$ and $\boldsymbol{r}_{i}(t)=[c_{2}(t),c_{4}(t),c_{6}(t)]$ of unit osculating invariant circle over eight revolutions, initialized at apolune. The norm of each vector field represents the projected length of the unit osculating invariant circle along each axis of the co-moving frame. Orange curves correspond to the first revolution, where each vector rotates by the rotation number $\varrho$.
(a) LVLH (RTN) frame
(b) TNW frame
@@
Moreover, applying SVD to the eigenvectors projected onto the two-dimensional co-moving frame allows for the evaluation of the minimum planar separation. Fig. 8 illustrates the evolution of the minimum separation in each plane of the LVLH (RTN) frame and the TNW frame along an absolute orbit, starting from the apolune. Only the TN plane has a nonzero separation for both frames, whereas the separations in the TW(RT) and RN(NW) planes collapse to zero at perilune (0.5 orbit). Furthermore, the separation in the TW plane increases by resolving the QPRIT in the TNW frame, whereas additional singularities are generated in the TN plane separation. This exemplifies the limitation of traditional safety measurements using the planar separations and motivates the evaluations in the local toroidal frame.
@@ 4 Passively Safe Reconfiguration using Local Toroidal Coordinates
Based on the above discussion, an OCP is constructed to provide a safe and optimal reconfiguration of spacecraft swarms. The discrete-time OCP in the nonsingular LTC is presented as follows:
$\min_{\{\boldsymbol{\zeta}_{\text{ns},k}^{(j)}\}_{k=0}^{N},\{\boldsymbol{u}_{k}^{(j)}\}_{k=0}^{N}}\sum_{j=1}^{N_{sc}}\mathcal{J}_{j}=\sum_{j=1}^{N_{sc}}\sum_{k=0}^{N}\|\boldsymbol{u}_{k}^{(j)}\|$ (17a)
s.t. $\boldsymbol{\zeta}_{\text{ns},0}^{(j)}=\boldsymbol{\zeta}_{\text{ns},i}^{(j)},\quad\boldsymbol{\zeta}_{\text{ns},f}^{(j)}=\boldsymbol{\zeta}_{\text{ns},N}^{(j)}+B_{N}\boldsymbol{u}_{N},$ $\forall j$ (17b)
@@
where $A_{k}=T_{k+1}^{-1}\Phi(t_{k+1},t_{k})T_{k},\quad B_{k}=\begin{bmatrix}\boldsymbol{0}_{3}\\ R^{-1}_{k}\end{bmatrix}.$ (17f)
The control input is an impulsive velocity change added to the $j$-th deputy within the TNW frame, denoted as $\boldsymbol{u}_{k}^{(j)}=\Delta\boldsymbol{v}_{k}^{(j)}$. The minimum-fuel problem is equivalent to the minimization of the sum of the $l_{2}$-norm of impulsive maneuvers. Using Eq. (13), the control input matrix $B_{k}$ is derived from the last three columns of $T_{k}^{-1}$. Eq. (17d) presents the passive safety constraint that is applied to two spacecraft $\boldsymbol{\zeta}_{\text{ns},k}^{(j)}$ and $\boldsymbol{\zeta}_{\text{ns},k}^{(j^{\prime})}$, where $j=0$ denotes the chief. Two passive safety formulations are discussed in the following subsections. Finally, Eq. (17e) prohibits a control near perilune because control around perilune is not desirable due to the high sensitivity, which leads to a large nonlinear error of the linearized dynamics. The no-control windows during transfers are heuristically defined in this paper.
Because the reference orbit may be highly eccentric (e.g., NRHO), a time regularization scheme is considered to distribute the discrete-time control nodes to reduce the nonlinear errors emanating from the dynamics linearization. The time discretization $dt$ is regularized to a pseudo-time discretization $d\tau$ by a scaling function of the chief’s position and velocity as $g(\boldsymbol{y})$ as $dt=g(\boldsymbol{y})d\tau\ \Leftrightarrow\ dt/d\tau=g(\boldsymbol{y})$. This leads to the reformulation of the dynamics in the pseudo-time domain as follows:
@@
A set of regularized time epochs can be determined by propagating the above dynamics over the pseudo-time interval $[\tau_{0},\tau_{f}]$, where the pseudo-time epochs are distributed uniformly. The retrieved time epochs that correspond to the pseudo-time epochs represent a vector of the regularized time epochs. In this paper, the Sundman transformation for the two-body system [43] is employed, considering that NRHOs are close to an elliptic orbit. The regularization function is defined as $g(\boldsymbol{y})=r^{\alpha}$, where $\alpha=1$ is a user-defined parameter.

### v2

@@ 3.2 Local Toroidal Coordinates
One can also define LTC by expressing the eigenspace component in polar coordinates (cf. Eq. (12)) as $\boldsymbol{\zeta}_{\text{geo}}:=[\varepsilon,\theta,h,\dot{\varepsilon},\dot{\theta},\dot{h}]^{\top}$, which is referred to as geometric LTC. This approximation holds for the relative motion around any periodic orbit with oscillatory mode thanks to the linearization adopted in Ref. [19], and the validity of its accuracy is directly affected by the size and direction of the invariant circle (i.e., eigenspace) at each phase of the reference orbit. A similar mapping from the geometric LTC to Cartesian coordinates is available [39], although the transformation is no longer a function of only the states on a reference orbit but also a time derivative of nonsingular LTC. Nonetheless, the geometric LTC provides a convenient characterization of an invariant circle’s size, thereby facilitating the specification of boundary conditions and waypoints for trajectory design.
The first-order condition to be on a QPRIT (QPRO) is to stay on the eigenspace spanned by $\boldsymbol{w}_{r}(t)$ and $\boldsymbol{w}_{i}(t)$. This leads to the condition $h=\dot{\alpha}=\dot{\beta}=\dot{h}=0$, or using geometric LTC, $h=\dot{\varepsilon}=\dot{\theta}=\dot{h}=0$ [40]. While other coordinates can still describe the deputy’s relative motion, the (first-order) quasi-periodicity linked to the oscillatory mode specific to the toroidal coordinate set cannot be maintained.
The quasi-periodic condition allows for arbitrary choices of the variables $(\alpha,\beta)$ or $(\varepsilon,\theta)$, defining both the size of the invariant circle and its phase. However, nonzero values in $(\dot{\alpha},\dot{\beta})$ or $(\dot{\varepsilon},\dot{\theta})$ result in a drift in the $h$-component, breaking the quasi-periodicity.
@@ 3.3 Safety evaluation of QROs
This subsection demonstrates the connection between the ROE and LTC based on the geometric configuration of the osculating relative orbit and the minimum separation. Although the approach is general, the application case of this section is on the Earth-Moon $L_{2}$ 9:2 South NRHO, which is the orbit of the Lunar Gateway [2, 55] and has an orbital period of 6.56 days.
3.3.1 Comparison between the LVLH and VNB
frame
The first-order approximation of a QPRIT resolved in the LVLH and VNB frames is illustrated in Figs. 5(a) and 5(b), respectively. Both figures depict the same QPRIT, where the fixed point is anchored at apolune with a scaling factor $\varepsilon=200$ m. The surfaces of the QPRIT are generated by the continuation of invariant circles (cf. Eq. (12)) along the reference orbit, shown as a gray surface. Additionally, fifty osculating invariant circles, sampled uniformly over one orbital period, are shown as blue ellipses. In the VNB frame, the semimajor axes of the invariant circles align with the chief’s velocity direction, producing a more compact structure. In contrast, for highly eccentric orbits like NRHOs, the angular offset of the QPRIT in the LVLH frame is evident. Previous works [40, 39, 37] report a QPRIT structure of NRHO characterized by two disk-like surfaces for a wide range of period, a pattern arising from positional displacement relative to the spacecraft velocity in the synodic frame. The reduced density of invariant circles within these disks reflects higher relative velocities and greater radial uncertainty.
@@ 3.3.2 Geometric Interpretation of Osculating Relative Invariant Circle
Two-dimensional projections of an osculating relative orbit in the PR2BP and an osculating invariant circle in the RMBPs are shown in Figs. 6(a) and 6(b), respectively. For a relative motion around a Keplerian orbit, a six-dimensional variable $\boldsymbol{\kappa}=\in\mathbb{R}^{6}$ parameterizes a relative orbit of the deputy with respect to the chief (cf. Eq. (10)), with additional independent variable $\rho(t)=1+e\cos\nu(t)$, where $e$ and $\nu(t)$ denote the eccentricity and true anomaly of the chief, respectively. In this paper, quasi-nonsingular ROE in near-circular orbits [5] and its extension to elliptic orbits [36] are chosen for $\boldsymbol{\kappa}$, which correspond to integration constants of the linearized equations of motion [3, 4]. In Fig. 6(a), $\kappa_{1}=0$ is enforced to realize a non-drifting bounded relative motion.
The position-components of the osculating invariant circle in Eq. (12) can be parameterized within the local frame as
[…]
$\tan\phi_{x}(t)=c_{1}(t)/c_{2}(t),\qquad\tan\phi_{y}(t)=c_{3}(t)/c_{4}(t),\qquad\tan\phi_{z}(t)=c_{5}(t)/c_{6}(t)$ (18)
$\boldsymbol{r}_{r}(t)=c_{1}(t)\hat{\boldsymbol{\imath}}+c_{3}(t)\hat{\boldsymbol{\jmath}}+c_{5}(t)\hat{\boldsymbol{k}},\qquad\boldsymbol{r}_{i}(t)=c_{2}(t)\hat{\boldsymbol{\imath}}+c_{4}(t)\hat{\boldsymbol{j}}+c_{6}(t)\hat{\boldsymbol{k}}$ (19)
@@
(b) Osculating invariant circle of a QPRO (RMBP) in the VNB
frame based on the parametrization in Eq. (17).
@@
Figure 6: Comparison of the geometric parametrization of osculating relative orbit in the PR2BP and the osculating invariant circle in the RMBPs.
@@
Vectors $c_{1}(t),c_{2}(t)]$, $c_{3}(t),c_{4}(t)]$, and $c_{5}(t),c_{6}(t)]$ provide the amplitudes of the projected ellipse in each plane as well as the oscillation phase. As stated in Sec. 3.1.1, $\boldsymbol{\kappa}$ is time-invariant [^5] When perturbations (e.g., spherical harmonics, drag, etc.) are added to the Keplerian dynamics, the osculating ROE also experiences a drift, although it does not induce any quasi-periodic motion [5, 6, 16, 36]. Nonetheless, the magnitude of the perturbation on the eigensystem is negligible (i.e., the eigenvectors of the monodromy matrix are almost 1) and can be treated as a secular drift in an averaged motion over multiple revolutions in a phase space. in the R2BP, allowing the explicit expression of the minimum planar separation in each projected plane [5]. In contrast, vector $\boldsymbol{c}(t)\in\mathbb{R}^{6}$ is time-varying in RMBPs, where a combined analysis of the magnitudes and phases of this vector is necessary to assess the passive safety of a QPRO.
@@
Fig. 7 presents the evolution of $\boldsymbol{c}(t)$ over eight revolutions, resolved in both the LVLH and VNB frames. The history is computed using a constant size $\varepsilon=1$ for the osculating invariant circle, ensuring preservation of the geometric structure. The norm of each vector field represents the projected length of the unit osculating invariant circle along each axis of the co-moving frame. Fig. 7 is analogous to the plots of the vector fields of relative eccentricity vector $\kappa_{3},\kappa_{4}]$ and relative inclination vector $\kappa_{5},\kappa_{6}]$ in the quasi-nonsingular ROE [5, 35] because the magnitude of each vector represents the amplitude of osculating invariant circle in each direction of the axis ($c_{12},c_{34}$, and $c_{56}$), and the magnitudes of the two vectors and the phase difference between them determine the minimum separation in the corresponding two-dimensional projected plane. However, note again that the vectors $c_{1}(t),c_{2}(t)]$, $c_{3}(t),c_{4}(t)]$, and $c_{5}(t),c_{6}(t)]$ for a unit invariant circle are determined by the local eigensystem, whereas the relative eccentricity and inclination vectors are free design variables of the deputy.
@@
From each vector field illustrated in Fig. 7, a quasi-periodic behavior is confirmed with a drastic change in its amplitude. Note that the equivalent vectors in relative motion for the PR2BP (i.e., ROE) would be stationary (for a near-circular orbit) or oscillatory about a line segment (for an elliptic orbit), with possible slow secular drifts due to perturbations. The evolution of the eigenvector can be characterized by two angles $\theta_{0},\theta_{1}]$ by separating the magnitude and the phase as $\boldsymbol{w}(t)=w_{r}(\theta_{0}(t))\hat{\boldsymbol{w}}_{r}(\theta_{1}(t))+w_{i}(\theta_{0}(t))\hat{\boldsymbol{w}}_{i}(\theta_{1}(t))i$. The first angle, $\theta_{0}$, determines the magnitude of the vector corresponding to the fixed point on the reference orbit. In contrast, the second angle, $\theta_{1}$, captures the secular drift of the vector over multiple orbital revolutions, reflecting the geometric structure of the two-dimensional QPIT and QPRIT (cf. Fig. 3(b)).
The angular rate of this secular rotation $\dot{\theta}_{1}$ remains constant over the three vector fields and corresponds to the angle of the complex eigenvalue of the oscillatory mode, known as rotation number $\varrho$ [28, 29]. For example, an eigenvalue associated with the oscillatory mode of the monodromy matrix of the 9:2 South NRHO is $\lambda\simeq 0.6845-0.7290i$. This leads to the rotation number of $\varrho\simeq-46.80^{\circ}$, which coincides with the angular displacement between the initial and terminal points of the orange segment in Fig. 7. For both Figs. 7(a) and 7(b), the minimum amplitude in the N-direction collapses to nearly zero, as shown in the history of $c_{3}(t),c_{4}(t)]$. A significant difference between Figs. 7(a) and 7(b) lies in the vector $c_{5}(t),c_{6}(t)]$, which represents the amplitude of the invariant circle in the R(W)-direction. This is illustrated by the contraction of the amplitude in the W-direction in the VNB frame compared to the R-direction in the LVLH frame.
@@
(b) VNB frame
Figure 7: Time history of the components of the eigenvectors $\boldsymbol{r}_{r}(t)$ and $\boldsymbol{r}_{i}(t)$ corresponding to the unit osculating invariant circle. Orange curves correspond to the first revolution, where each vector rotates by the rotation number $\varrho$.
(a) LVLH frame
(b) VNB frame
@@
Previous studies provide analysis on the minimum positional separation by performing an SVD [40] or modal decomposition [24]. Applying SVD to the eigenvectors projected onto the two-dimensional co-moving frame allows for the evaluation of the minimum planar separation, which offers additional insights into the design of the safety region. Fig. 8 illustrates the evolution of the minimum separation in each plane of the LVLH frame and the VNB frame along the full orbit, starting from the apolune. Only the ZX/BV-plane has a nonzero separation for both frames, whereas the separations in the XY/VN- and YZ/NB-planes collapse to zero at perilune (0.5 orbit). Furthermore, the separation in the NB plane increases by resolving the QPRIT in the VNB frame, whereas additional singularities are generated in the VN- plane separation. Note that although the separation increases with a larger QPIT radius ($\varepsilon$), this trend is specific to the orbit. In the PR2BPs, an appropriate relative orbit configuration can always secure planar separation in the YZ plane, which is known as relative eccentricity/inclination-vector separation [7, 5]. This also motivates a unique passive safety formulation for the relative motion in the RMBPs.
@@ 4 Passively Safe Reconfiguration using Local Toroidal Coordinates
This section presents a passively safe, fuel-optimal swarm reconfiguration method in cislunar relative motion. Assuming that the chief is on a reference orbit, the focus is on transferring an $N_{sc}$ deputy spacecraft between two passively safe QPROs, where the initial and terminal conditions are defined by $\boldsymbol{\zeta}_{\text{ns},i}^{(j)}$ and $\boldsymbol{\zeta}_{\text{ns},f}^{(j)}$ for each $j=1,\dots,N_{sc}$, with a fixed transfer time $t_{f}$ while ensuring passive safety throughout the transfer. Such safe transfers are crucial for missions involving on-orbit inspection or formation flying.
To facilitate numerical optimization and enable efficient trajectory planning under mission constraints, the problem is cast in a discrete-time framework. This approach also aligns with the use of linearized dynamics models and impulsive control inputs. With a sufficiently fine time discretization, this problem can be formulated as a discrete-time OCP using the linearized dynamics on the nonsingular LTC within a local frame as follows:
$\min_{\{\boldsymbol{\zeta}_{\text{ns},k}^{(j)}\}_{k=0}^{N},\{\boldsymbol{u}_{k}^{(j)}\}_{k=0}^{N}}\$$\sum_{j=1}^{N_{sc}}\mathcal{J}_{j}=\sum_{j=1}^{N_{sc}}\sum_{k=0}^{N}\|\boldsymbol{u}_{k}^{(j)}\|$ (20a)
subject to $\boldsymbol{\zeta}_{\text{ns},i}^{(j)}=\boldsymbol{\zeta}_{\text{ns},0}^{(j)},\qquad\boldsymbol{\zeta}_{\text{ns},f}^{(j)}=\boldsymbol{\zeta}_{\text{ns},N}^{(j)}+B_{N}\boldsymbol{u}_{N},$ $\forall j$ (20b)
@@
where $A_{k}=T_{k+1}^{-1}\Phi(t_{k+1},t_{k})T_{k},\qquad B_{k}=\begin{bmatrix}\boldsymbol{0}_{3}\\ R^{-1}_{k}\end{bmatrix}$. (20f)
The control input $\boldsymbol{u}_{k}^{(j)}$ is an impulsive velocity change added to the $j$-th deputy within the VNB frame. The minimum-fuel problem is equivalent to the minimization of the sum of the $l_{2}$-norm of impulsive maneuvers. Using Eq. (14), the control input matrix $B_{k}$ is derived from the last three columns of $T_{k}^{-1}$. Eq. (20d) presents the passive safety constraint that is applied to two spacecraft $\boldsymbol{\zeta}_{\text{ns},k}^{(j)}$ and $\boldsymbol{\zeta}_{\text{ns},k}^{(j^{\prime})}$, where $j=0$ denotes the chief. Two passive safety formulations are discussed in the following subsections.
To avoid control near perilune, where system sensitivity is high and the linearized dynamics incur substantial nonlinear errors, Eq. (20e) enforces a no-control condition. These no-control windows, denoted by the heuristic intervals $k\in[k_{p}^{-},k_{p}^{+}]$, define time segments during which control actions are not allowed.
Because the reference orbit may be highly eccentric (e.g., NRHO), a time regularization scheme is considered to distribute the discrete-time control nodes to reduce the nonlinear errors emanating from the dynamics linearization. The time discretization $dt$ is regularized to a pseudo-time discretization $d\tau$ by a scaling function of the chief’s position and velocity as $g(\boldsymbol{y})$ as $dt=g(\boldsymbol{y})d\tau\ \Leftrightarrow\ dt/d\tau=g(\boldsymbol{y})$. This leads to the reformulation of the dynamics in the pseudo-time domain as
@@
A set of regularized time epochs can be determined by propagating the above dynamics over the pseudo-time interval $\tau_{0},\tau_{f}]$, where the pseudo-time epochs are distributed uniformly. The retrieved time epochs that correspond to the pseudo-time epochs represent a vector of the regularized time epochs. In this paper, the Sundman transformation for the two-body system [56] with respect to the Moon is employed, considering that NRHOs are close to an elliptic orbit. The regularization function is defined as $g(\boldsymbol{y})=r^{\alpha}$, where $\alpha=1$ is a user-defined parameter.